THE CYCLIC CO_{2} TRIMER: OBSERVATION OF TWO PARALLEL BANDS AND DETERMINATION OF INTERMOLECULAR OUT-OF-PLANE TORSIONAL FREQUENCIES

\underline{M. AFSHARI}, M. DEHGHANY, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, 2500 University Drive NW, Calgary, Alberta T2N 1N4, Canada; A.R.W. MCKELLAR, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

Previously, two distinct isomers of the CO_{2} trimer have been identified by means of infrared spectroscopy. The first is a symmetric top with a cyclic planar structure and the second is an asymmetric top barrel-shaped structure with C_{2} symmetry. We have observed two new parallel (ΔK = 0) bands of the cyclic CO_{2} trimer at 2364 and 2370.5 cm^{−1} which are each assigned as a combination of an intramolecular CO_{2} monomer ν_{3} stretch and an intermolecular out-of-plane torsion, giving the torsional frequencies of 12.3 and 18.8 cm^{−1}, respectively. The band at 2364 cm^{−1} is surprisingly strong and completely unperturbed, providing a rare and near perfect example for a parallel band of a symmetric top molecule with C_{3h} symmetry and zero nuclear spins. The second band at 2370.5 cm^{−1} is somewhat weaker and has not been fully analysed because of spotty coverage of our diode lasers. However, it is still easily assigned as the second anticipated combination band involving an out-of-plane torsional mode of the cyclic CO_{2} trimer. The trimers are generated in a pulsed supersonic expansion from a slit-jet nozzle and probed with a tunable infrared diode laser.